# SEACrowd: A Multilingual Multimodal Data Hub and Benchmark Suite for Southeast Asian Languages

### v1

@@ Abstract
Southeast Asia (SEA) is a region rich in linguistic diversity and cultural variety, with over 1,300 indigenous languages and a population of 671 million people. However, prevailing AI models suffer from a significant lack of representation of texts, images, and audio datasets from SEA, compromising the quality of AI models for SEA languages. Evaluating models for SEA languages is challenging due to the scarcity of high-quality datasets, compounded by the dominance of English training data, raising concerns about potential cultural misrepresentation. To address these challenges, we introduce SEACrowd, a collaborative initiative that consolidates a comprehensive resource hub [^1] https://seacrowd.github.io/seacrowd-catalogue/ that fills the resource gap by providing standardized corpora [^2] https://github.com/SEACrowd/seacrowd-datahub/ in nearly 1,000 SEA languages across three modalities. Through our SEACrowd benchmarks, we assess the quality of AI models on 36 indigenous languages across 13 tasks, offering valuable insights into the current AI landscape in SEA. Furthermore, we propose strategies to facilitate greater AI advancements, maximizing potential utility and resource equity for the future of AI in SEA.
@@
Holy Lovenia★,1,2 Rahmad Mahendra★,3,2 Salsabil Maulana Akbar★,2 Lester James V. Miranda★,4 Jennifer Santoso★,5Elyanah Aco★,6Akhdan Fadhilah★,7Jonibek Mansurov★,8Joseph Marvin Imperial★,9,10 Onno P. Kampman★,11Joel Ruben Antony Moniz★,6Muhammad Ravi Shulthan Habibi★,3,2Frederikus Hudi★,12,13 Railey Montalan★,1Ryan Ignatius6Joanito Agili Lopo14William Nixon15Börje F. Karlsson16James Jaya6 Ryandito Diandaru6Yuze Gao6Patrick Amadeus15Bin Wang6Jan Christian Blaise Cruz17Chenxi Whitehouse18 Ivan Halim Parmonangan19Maria Khelli15Wenyu Zhang6Lucky Susanto20Reynard Adha Ryanda21 Sonny Lazuardi Hermawan22Dan John Velasco17Muhammad Dehan Al Kautsar15Willy Fitra Hendria6 Yasmin Moslem23Noah Flynn24Muhammad Farid Adilazuarda8Haochen Li6Johanes Lee15R. Damanhuri25 Shuo Sun6Muhammad Reza Qorib26Amirbek Djanibekov8Wei Qi Leong1Quyet V. Do27Niklas Muennighoff28 Tanrada Pansuwan18Ilham Firdausi Putra6Yan Xu29,27Ngee Chia Tai1Ayu Purwarianti6,30 Sebastian Ruder31William Tjhi1Peerat Limkonchotiwat★,32Alham Fikri Aji★,8Sedrick Keh★,33 Genta Indra Winata★,2Ruochen Zhang★,34Fajri Koto★,8,2Zheng-Xin Yong★,34Samuel Cahyawijaya★,27,2 1AI Singapore 2IndoNLP 3Universitas Indonesia 4Allen Institute for Artificial Intelligence 5RevComm, Inc. 6Independent Researcher 7Tohoku University 8MBZUAI 9University of Bath 10National University Philippines 11MOH Office for Healthcare Transformation (MOHT) 12NAIST 13Works Applications Lab 14Universitas Gadjah Mada 15Institut Teknologi Bandung 16Beijing Academy of Artificial Intelligence (BAAI) 17Samsung Research Philippines 18University of Cambridge 19Queensland University of Technology 20Monash University Indonesia 21Imperial College London 22Independent Design Engineer 23Bering Lab 24Amazon 25Universitas Diponegoro 26NUS 27HKUST 28Contextual AI 29Huawei Noah’s Ark Lab 30Prosa.ai 31Cohere 32VISTEC 33Toyota Research Institute 34Brown University Major contributors
@@ 1 Introduction
Despite the Southeast Asia (SEA) region being home to 1,300 indigenous languages (18% of the world’s languages) and 671 million people (8.75% of the world’s population), the representation of texts, images, and audio datasets from this region is significantly lacking in machine learning pre-training models. This deficiency negatively impacts the model quality for SEA languages. The language coverage of SEA languages in two common pre-training resources, Common Crawl [^3] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages and C4 Xue et al. (2021), is extremely limited, with only 2.36% (in 11 languages) and 10.62% (in 11 languages), respectively. In modalities beyond text, the representation is even more limited. For instance, Common Voice, one of the largest multilingual speech corpora, includes only 6 SEA indigenous languages Conneau et al. (2021); Ardila et al. (2020), and LAION-5B, one of the largest multilingual vision-language (VL) corpora, includes only 12 SEA indigenous languages Schuhmann et al. (2022). While datasets for other SEA indigenous languages may exist, they are often scattered, insufficiently documented, or varied in quality and formatting, thereby making access and usage challenging Cahyawijaya et al. (2023a); Joshi et al. (2020); Aji et al. (2023).
In terms of evaluation, the sparse availability of high-quality test tasks for these languages also complicates evaluating models for SEA languages. Despite there being 1,300+ languages in the SEA region, prior works  Winata et al. (2023); Cahyawijaya et al. (2021); Koto and Koto (2020); Zhang et al. (2024); Wang et al. (2024); Nguyen et al. (2023); Leong et al. (2023); Yong et al. (2023) have only evaluated fewer than 10 SEA languages collectively. The actual performance of current models on most SEA languages remains largely unknown.
Moreover, the dominance of Anglocentric training data potentially results in cultural bias when generating texts, images, or audio in underrepresented SEA languages Søgaard (2022); Talat et al. (2022). Further, Durmus et al. (2023); AlKhamissi et al. (2024); Cahyawijaya et al. (2024a) have shown that the learned representations in LMs often fail to reflect local cultural values in SEA Koto et al. (2024); Liu et al. (2024); Adilazuarda et al. (2024). This raises concerns about the ability of current LLMs to generate natural, high-quality texts for this region. Furthermore, the discrepancy in language support creates language barriers in technological access and risks marginalizing minority groups who do not speak the dominant language.
@@
  • We bridge the resource gap by centralizing and standardizing $\sim$500 corpora in nearly 1,000 SEA languages in SEACrowd, a comprehensive and standardized resource center, across 3 modalities: text, image, and audio.
  • We close the evaluation gap in SEA languages with the SEACrowd Benchmarks, which cover 38 SEA indigenous languages on 13 tasks across 3 modalities, providing insights into the performance of a diverse spectrum of AI models. Further, our study reveals that the generative outputs of existing LLMs exhibit a closer resemblance to translationese rather than natural data in 9 SEA languages.
  • We offer insights and strategies for the future development of AI in SEA, aiming to promote a sustainable and prosperous future through continuous AI advancements.
@@ 2 SEACrowd
We invited contributors to submit datasheet forms for publicly available datasets across all modalities including text, audio, and image in SEA languages and/or cultures. These datasheets include detailed information about each dataset, such as data subset(s), description, task, language, license, URL access, annotation method(s), annotation validation, relevant publications, publication venue, and data splits. For each submission, we manually verify and correct it as necessary to ensure datasheet accuracy.
@@ 3.2 Baselines
To evaluate the zero-shot performance of instruction-tuned LLMs on SEA languages, we benchmark two commercial, i.e., GPT-4 OpenAI et al. (2024) and Command-R [^8] https://docs.cohere.com/docs/command-r, and 17 open-source baselines, the majority of which are $\sim$7B-13B parameters. We categorize the open-source baselines according to the language(s) coverage in pre-training and/or instruction tuning, i.e., 1) English: Llama3 Touvron et al. (2023), Mistral Jiang et al. (2023), and Falcon Almazrouei et al. (2023); 2) Multilingual: AYA-101, AYA-23 Üstün et al. (2024), mT0, BLOOMZ Muennighoff et al. (2022), and BactrianX-Llama Li et al. (2023a); 3) SEA regional: SEA-LION Singapore (2023), Sailor Dou et al. (2024), and SeaLLM Nguyen et al. (2023); and 4) SEA country-specific: Cendol-mT5, Cendol-Llama2 Cahyawijaya et al. (2024b), Merak Ichsan (2023) from Indonesia, WangchanX-Llama3 Phatthiyaphaibun et al. (2024) from Thailand, and Malaysian-Llama3999https://huggingface.co/mesolitica/malaysian-llama-3-8b-instruct-16k from Malaysia.
@@
We evaluate the zero-shot performances of state-of-the-art multilingual pre-trained speech models in transcribing speech in SEA languages. Specifically, we consider Whisper v3 Radford et al. (2023), MMS 1B Pratap et al. (2024), and Seamless M4T v2 Communication et al. (2023), which have shown proficiency in accurately transcribing multiple languages without fine-tuning. Additionally, we include models that are fine-tuned on specific language(s), SEA or English, based on 1) Wav2Vec2 XLSR Conneau et al. (2021) and 2) XLS-R Babu et al. (2021), known for their cross-lingual speech representation learning by pre-training on raw speech waveforms across diverse languages, with XLS-R offering broader language coverage, and 3) Whisper, which leverages weakly supervised pre-training on spectrograms of speech in diverse languages. The specific fine-tuned models are evaluated: XLSR on ind, jav, sun; XLSR and Whisper on Indonesian (ind); XLSR and Whisper on Thai (tha); XLS-R on Tagalog (tgl); XLS-R on Burmese (mya); XLS-R and Whisper on Khmer (khm); and XLSR on English (eng). See Appendix G.2 for details.
@@ 4.1 State-of-the-Art Models on SEA languages
Figure 2(a) and 2(b) illustrate the overall model performance of the LLM baselines in SEA languages for both NLU tasks and NLG tasks. In our NLU evaluation, AYA-101, a large multilingual instruction-tuned language model covering 101 languages, demonstrates the best zero-shot performance. It is followed by the commercial baselines, which achieve a median of $\sim$0.6 weighted F1-score. Sailor and SeaLLM, models specifically trained with SEA languages, also display competitive performance. Similarly, mT0 exhibits strong generalization abilities due to its exposure to $\sim$100 languages in pre-training, including those from the SEA region Muennighoff et al. (2022). In contrast, most English and SEA country-specific baselines perform less effectively, likely due to their narrow focus on English or a limited set of SEA languages, such as Indonesian languages for Cendol and Thai for WangchanX-Llama3. Similar and consistent trends are observed for the NLG tasks, except for GPT-4’s poorer performance across summarization, abstractive or extractive QA, and MT, indicating its ineffectiveness in producing acceptable outputs in SEA languages for these tasks. Appendix G.4 describes the performance of LLMs per language.
@@
Figure 4: Existing VLMs produce subpar image captions in SEA languages.
@@
Figure 3 presents the off-the-shelf speech model performance on ASR across languages in SEA, measured by the error rate percentage. 9 of the 15 SEA languages in our speech evaluation belong to the Austronesian language family. The other 6 are khm and vie, which belong to Austro-Asiatic, cnh and mya belong to Sino-Tibetan, as well as tha and vie belong Kra-Dai language family. Generally, the multilingual pre-trained baselines have a competitive generalization capability across the languages, although it varies by language. For instance, Whisper v3 demonstrates significantly higher effectiveness for national languages such as ind, zlm, fil, tha, and vie, while performing less optimally for other indigenous languages. Conversely, Seamless M4T v2 shows a more balanced performance across the languages. Regarding fine-tuned baselines, error rates decrease for their seen languages. The fine-tuned Whisper models, however, manage to better optimize for the target language while retaining their original capabilities in other SEA languages compared to their Wav2Vec2 XLSR and XLS-R counterparts, despite both having been pre-trained in a multilingual manner. This observation aligns with the findings of Rouditchenko et al. (2023), who find that the number of hours seen per language and language family during pre-training is predictive of how the models compare, in which Whisper’s pre-training data duration for these four language families exceeds that of XLSR.
@@
However, when prompted in eng, the performance of these multilingual baselines varies notably. PaliGemma’s performance diminishes completely, while mBLIP’s performance shows both increases and decreases across different SEA languages. This raises the question of whether the multilingual VLMs can maintain consistent performance across different languages used in the instructions and the tasks. It highlights the need for further research into the mechanisms that drive these variations and how to achieve robust multilingual performance in VLMs across diverse linguistic contexts. Understanding these dynamics is crucial for improving VLMs’ generalization capabilities and ensuring equitable performance across all languages, despite most related works focusing on monolingual visual instruction tuning Liu et al. (2023b); Gong et al. (2023); Zhu et al. (2024).
4.2 Generation Quality in SEA Languages: Translationese vs. Natural
@@
To analyze the generation quality of LLMs in SEA languages, we build a text classifier to discriminate between translationese and natural texts Riley et al. (2020). We construct a translationese classification training and testing dataset using 49 and 62 data subsets, respectively, covering approximately 39.9k and 51.5k sentences across 9 SEA languages: English (eng), Indonesian (ind), Khmer (khm), Lao (lao), Burmese (mya), Filipino (fil), Thai (tha), Vietnamese (vie), and Malay (zlm). The training and test data are detailed in Appendix H.1.
@@
We evaluate the generation quality of LLMs in 9 SEA languages by generating answers to natural, general, and safety questions from Sea-Bench Nguyen et al. (2023). As shown in Table 2(a), LLMs with extensive language coverage but less focus on SEA languages, e.g., AYA-101 Üstün et al. (2024), GPT-4 OpenAI et al. (2024), mT0 Muennighoff et al. (2023); Xue et al. (2021), and Llama3 AI@Meta (2024), tend to produce natural sentences less than 20% of the time. In contrast, models with narrower language coverage but a greater focus on SEA languages, such as Cendol-Llama2 Cahyawijaya et al. (2024b), Sailor Dou et al. (2024), AYA-23 Aryabumi et al. (2024), and SEA-LION Singapore (2023), generate more natural sentences over 35% of the time.

### v4

@@ Abstract
Southeast Asia (SEA) is a region characterized by rich linguistic diversity and cultural variety, with over 1,300 indigenous languages and a population of 671 million people. However, the performance of contemporary AI models for SEA languages is compromised by a significant lack of representation of texts, images, and auditory datasets from SEA. Evaluating models for SEA languages is challenging due to the scarcity of high-quality datasets, compounded by the predominance of English training data, which raises concerns regarding potential cultural misrepresentation. To address these challenges, we introduce SEACrowd, a collaborative initiative that consolidates a comprehensive resource hub [^1] https://seacrowd.github.io/seacrowd-catalogue/ to bridge the resource gap by providing standardized corpora and benchmarks [^2] https://github.com/SEACrowd/seacrowd-datahub/ in nearly 1,000 SEA languages across three modalities. We assess the performance of AI models on 36 indigenous languages across 13 tasks included in SEACrowd, offering valuable insights into the current AI landscape in SEA. Furthermore, we propose strategies to facilitate greater AI advancements, maximizing potential utility and resource equity for the future of AI in Southeast Asia.
@@
Holy Lovenia★,1,2 Rahmad Mahendra★,3,2 Salsabil Maulana Akbar★,2 Lester James V. Miranda★,4 Jennifer Santoso★,5Elyanah Aco★,6Akhdan Fadhilah★,7Jonibek Mansurov★,8Joseph Marvin Imperial★,9,10 Onno P. Kampman★,11Joel Ruben Antony Moniz★,6Muhammad Ravi Shulthan Habibi★,3,2Frederikus Hudi★,12,13 Railey Montalan★,1Ryan Ignatius6Joanito Agili Lopo14William Nixon15Börje F. Karlsson16James Jaya6 Ryandito Diandaru6Yuze Gao6Patrick Amadeus15Bin Wang6Jan Christian Blaise Cruz8,17Chenxi Whitehouse18 Ivan Halim Parmonangan19Maria Khelli15Wenyu Zhang6Lucky Susanto20Reynard Adha Ryanda21 Sonny Lazuardi Hermawan22Dan John Velasco17Muhammad Dehan Al Kautsar15Willy Fitra Hendria6 Yasmin Moslem23Noah Flynn24Muhammad Farid Adilazuarda8Haochen Li6Johanes Lee15R. Damanhuri25 Shuo Sun6Muhammad Reza Qorib26Amirbek Djanibekov8Wei Qi Leong1Quyet V. Do27Niklas Muennighoff28 Tanrada Pansuwan18Ilham Firdausi Putra6Yan Xu29,27Ngee Chia Tai1Ayu Purwarianti6,30 Sebastian Ruder31William Tjhi1Peerat Limkonchotiwat★,32Alham Fikri Aji★,8Sedrick Keh★,33 Genta Indra Winata★,35,2Ruochen Zhang★,34Fajri Koto★,8,2Zheng-Xin Yong★,34Samuel Cahyawijaya★,31,27,2 1AI Singapore 2IndoNLP 3Universitas Indonesia 4Allen Institute for Artificial Intelligence 5RevComm, Inc. 6Independent Researcher 7Tohoku University 8MBZUAI 9University of Bath 10National University Philippines 11MOH Office for Healthcare Transformation (MOHT) 12NAIST 13Works Applications Lab 14Universitas Gadjah Mada 15Institut Teknologi Bandung 16Beijing Academy of Artificial Intelligence (BAAI) 17Samsung Research Philippines 18University of Cambridge 19Queensland University of Technology 20Monash University Indonesia 21Imperial College London 22Independent Design Engineer 23Bering Lab 24Amazon 25Universitas Diponegoro 26NUS 27HKUST 28Contextual AI 29Huawei Noah’s Ark Lab 30Prosa.ai 31Cohere 32VISTEC 33Toyota Research Institute 34Brown University 35Capital One Major contributors
@@ 1 Introduction
Despite Southeast Asia (SEA) being home to 1,300 indigenous languages (18% of the world’s languages) and 671 million people (8.75% of the world’s population), the representation of texts, images, and audio datasets from this region is significantly lacking in machine learning models. This deficiency adversely affects the model quality for SEA languages. The language coverage of SEA languages in two common pre-training resources, Common Crawl [^3] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages and C4 Xue et al. (2021), is extremely limited, with only 2.36% (in 11 languages) and 10.62% (in 11 languages), respectively. In modalities beyond text, the representation is even more limited. For instance, Common Voice, one of the largest multilingual speech corpora, includes six SEA indigenous languages Conneau et al. (2021); Ardila et al. (2020), and LAION-5B, one of the largest multilingual vision-language (VL) corpora, includes 12 SEA indigenous languages Schuhmann et al. (2022). Datasets for other SEA indigenous languages exist, but are often scattered, insufficiently documented, or varied in quality and formatting, thereby making access and usage challenging Cahyawijaya et al. (2023a); Joshi et al. (2020); Aji et al. (2023).
In terms of evaluation, the sparse availability of high-quality test sets for these languages also complicates evaluating models for SEA languages. Despite there being 1,300+ languages in the SEA region, prior works  Winata et al. (2023); Cahyawijaya et al. (2021); Koto and Koto (2020); Zhang et al. (2024); Wang et al. (2024); Nguyen et al. (2023); Leong et al. (2023); Yong et al. (2023) have only evaluated fewer than 10 SEA languages collectively. The actual performance of current models on most SEA languages remains largely unknown.
Moreover, the dominance of Anglocentric training data may result in cultural bias when generating texts, images, or audio in underrepresented SEA languages Søgaard (2022); Talat et al. (2022). Further, Durmus et al. (2023); AlKhamissi et al. (2024); Cahyawijaya et al. (2024a) have shown that the learned representations in large language models (LLMs) often fail to reflect local cultural values in SEA Koto et al. (2024); Liu et al. (2024); Adilazuarda et al. (2024). This raises concerns about the ability of current LLMs to generate natural, high-quality texts for this region. In addition, the discrepancy in language support creates language barriers in technological access and risks marginalizing minority groups who do not speak the dominant language.
@@
  • We bridge the resource gap by centralizing and standardizing $\sim$500 corpora in nearly 1,000 SEA languages in SEACrowd, a comprehensive and standardized resource center, across three modalities: text, image, and audio.
  • We close the evaluation gap in SEA languages with the SEACrowd Benchmarks, which cover 38 SEA indigenous languages on 13 tasks across 3 modalities, providing insights into the performance of a diverse spectrum of AI models. Further, our study reveals that the generative outputs of existing LLMs exhibit a closer resemblance to “translationese” rather than natural data in nine SEA languages.
  • We offer insights and strategies for the future development of AI in SEA.
@@ 2 SEACrowd
We invited contributors to submit datasheet forms Gebru et al. (2021) for publicly available datasets across all modalities including text, audio, and image in SEA languages and/or cultures. These datasheets include detailed information about each dataset, such as data subset(s), description, task, language, license, URL access, annotation method(s), annotation validation, relevant publications, publication venue, and data splits. For each submission, we manually verify and correct it as necessary to ensure datasheet accuracy.
@@ 3.2 Baselines
To evaluate the zero-shot performance of instruction-tuned LLMs on SEA languages, we benchmark two commercial, i.e., GPT-4 OpenAI et al. (2024) and Command-R [^8] https://docs.cohere.com/docs/command-r, and 17 open-source baselines, the majority of which are $\sim$7B-13B parameters. We categorize the open-source baselines according to the language(s) coverage in pre-training and/or instruction tuning, i.e., 1) English: Llama3 Touvron et al. (2023), Mistral Jiang et al. (2023), and Falcon Almazrouei et al. (2023); 2) Multilingual: AYA-101, AYA-23 Üstün et al. (2024), mT0, BLOOMZ Muennighoff et al. (2022), and BactrianX-Llama Li et al. (2023a); 3) SEA regional: SEA-LION Singapore (2023), Sailor Dou et al. (2024), and SeaLLM Nguyen et al. (2023); and 4) SEA country-specific: Cendol-mT5, Cendol-Llama2 Cahyawijaya et al. (2024b), and Merak Ichsan (2023) from Indonesia, WangchanX-Llama3 Phatthiyaphaibun et al. (2024) from Thailand, and Malaysian-Llama3999https://huggingface.co/mesolitica/malaysian-llama-3-8b-instruct-16k from Malaysia.
@@
We evaluate the zero-shot performance of state-of-the-art multilingual pre-trained speech models in transcribing speech in SEA languages. Specifically, we consider Whisper v3 Radford et al. (2023), MMS 1B Pratap et al. (2024), and Seamless M4T v2 Communication et al. (2023), which have shown proficiency in accurately transcribing multiple languages without fine-tuning. Additionally, we include models that are fine-tuned on specific language(s), SEA or English, based on 1) Wav2Vec2 XLSR Conneau et al. (2021) and 2) XLS-R Babu et al. (2021), known for their cross-lingual speech representation learning by pre-training on raw speech waveforms across diverse languages, with XLS-R offering broader language coverage, and 3) Whisper, which leverages weakly supervised pre-training on spectrograms of speech in diverse languages. The specific fine-tuned models are evaluated: XLSR on ind, jav, sun; XLSR and Whisper on Indonesian (ind); XLSR and Whisper on Thai (tha); XLS-R on Tagalog (tgl); XLS-R on Burmese (mya); XLS-R and Whisper on Khmer (khm); and XLSR on English (eng). See Appendix G.2 for details.
@@ 4.1 State-of-the-Art Models on SEA languages
Figure 2(a) and 2(b) illustrate the overall model performance of the LLM baselines in SEA languages for both NLU tasks and NLG tasks. In our NLU evaluation, AYA-101, a large multilingual instruction-tuned language model covering 101 languages, demonstrates the best zero-shot performance. It is followed by the commercial baselines, which achieve a median of $\sim$0.6 weighted F1-score. Sailor and SeaLLM, models specifically trained with SEA languages, also display competitive performance. Similarly, mT0 exhibits strong generalization abilities due to its exposure to $\sim$100 languages in pre-training, including those from the SEA region Muennighoff et al. (2022). In contrast, most English and SEA country-specific baselines perform less effectively, likely due to their narrow focus on English or a limited set of SEA languages, such as Indonesian languages for Cendol and Thai for WangchanX-Llama3. Similar and consistent trends are observed on MT task, while the baselines’ poorer scores on abstractive/extractive QA and summarization indicate their ineffectiveness in producing acceptable outputs in SEA languages for these tasks, which is especially pronounced in the open-source baselines. Appendix G.4 describes the performance of LLMs per language.
@@
Figure 4: Existing VLMs produce subpar image captions in SEA languages. We report CIDEr Vedantam et al. (2015).
@@
Figure 3 presents the off-the-shelf speech model performance on ASR across languages in SEA, measured by the error rate percentage. 9 of the 15 SEA languages in our speech evaluation belong to the Austronesian language family. The other 6 are khm and vie, which belong to Austro-Asiatic, cnh and mya belong to Sino-Tibetan, and tha and vie belong to the Kra-Dai language family. The multilingual pre-trained baselines have a competitive generalization capability across languages, although it varies by language. For instance, Whisper v3 demonstrates significantly higher effectiveness for national languages such as ind, zlm, fil, tha, and vie, while performing less optimally for other indigenous languages. Conversely, Seamless M4T v2 shows a more balanced performance across the languages. Regarding fine-tuned baselines, error rates decrease for their seen languages. The fine-tuned Whisper models, however, manage to better optimize for the target language while retaining their original capabilities in other SEA languages compared to their Wav2Vec2 XLSR and XLS-R counterparts, despite both having been pre-trained in a multilingual manner. This observation aligns with the findings of Rouditchenko et al. (2023), who find that the number of hours seen per language and language family during pre-training is predictive of how the models compare, in which Whisper’s pre-training data duration for these four language families exceeds that of XLSR.
@@
However, when prompted in eng, the performance of these multilingual baselines varies notably. PaliGemma’s performance collapses completely, while mBLIP’s performance shows both increases and decreases across different SEA languages. This raises the question of whether the multilingual VLMs can maintain consistent performance across different languages used in the instructions and the tasks. It highlights the need for further research into the mechanisms that drive these variations and how to achieve robust multilingual performance in VLMs across diverse linguistic contexts. Understanding these dynamics is crucial for improving VLMs’ generalization capabilities and ensuring equitable performance across all languages, despite most related works focusing on monolingual visual instruction tuning Liu et al. (2023b); Gong et al. (2023); Zhu et al. (2024).
4.2 Generation Quality in SEA Languages: Translationese vs. Natural Language
@@
To analyze the generation quality of LLMs in SEA languages, we build a text classifier to discriminate between translationese and natural texts Riley et al. (2020). We construct a translationese classification training and testing dataset using 49 and 62 data subsets, respectively, covering approximately 39.9k and 51.5k sentences across English (eng) and 8 SEA languages: Indonesian (ind), Khmer (khm), Lao (lao), Burmese (mya), Filipino (fil), Thai (tha), Vietnamese (vie), and Malay (zlm). The training and test data are detailed in Appendix H.1.
@@
We evaluate the generation quality of LLMs in 9 SEA languages by generating answers to natural, general, and safety questions from Sea-Bench Nguyen et al. (2023). As shown in Table 2(a), LLMs with extensive language coverage but less focus on SEA languages, e.g., AYA-101 Üstün et al. (2024), GPT-4 OpenAI et al. (2024), mT0 Muennighoff et al. (2023); Xue et al. (2021), and Llama3 AI@Meta (2024), tend to produce natural sentences less than 20% of the time. In contrast, models with narrower language coverage but a greater focus on SEA languages, such as Cendol-Llama2 Cahyawijaya et al. (2024b), Sailor Dou et al. (2024), AYA-23 Aryabumi et al. (2024), and SEA-LION Singapore (2023), generate natural sentences over 35% of the time.